We are CLOSED

Sensor Status Last Update
Open Switch Flipped OFF Aug 22, 2017 10:02 PM PDT
Front Door Closed Aug 22, 2017 11:07 PM PDT
Main Door Closed Aug 22, 2017 11:07 PM PDT
Office Motion No Movement since Aug 22, 2017 11:07 PM PDT
Shop Motion No Movement since Aug 22, 2017 11:07 PM PDT
Temperature 26.75 °C/ 80.15 °F Aug 23, 2017 6:30 AM PDT
Page Loaded at Aug 23, 2017 6:33 AM PDT